OpenStudy (anonymous):

through(3,-1) perp to y=8/3x+3 slope intercept form is it y=-3/8+1/8

OpenStudy (anonymous):

y=-3/8x+17/8

OpenStudy (anonymous):

Yes except you need an x after the -3/8... good job. -1 = (-3/8)*3 + b -1 = -9/8 + b -1 = -1 1/8 + b 1/8 = b You were write michaela.. just put the x after the -3/8

OpenStudy (anonymous):

1=-9/8 + b so 8/8+9/8 = 17/8

OpenStudy (anonymous):

btm y value is -1

OpenStudy (anonymous):

sorry I used 1...just saw it ; (

OpenStudy (anonymous):

getting too EXCITED about this....first night, lol!

OpenStudy (anonymous):

lol aw well thanks both of u!